40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed May 15 00:40:19 2019, gautam, Update, SUS, ETMY diagnosis plan ETMY_sensorSpectra.pdf
    Reply  Wed May 15 10:57:57 2019, gautam, Update, SUS, EY chamber opened 
       Reply  Wed May 15 17:46:24 2019, gautam, Update, SUS, ETMY inspection firstContactFiber.JPGETMY_sensorSpectra.pdf
          Reply  Wed May 15 19:36:29 2019, Koji, Update, SUS, ETMY instepction 
             Reply  Thu May 16 13:07:14 2019, gautam, Update, SUS, First contact residue removal etmy_sensors.pdfetmy_BRmode.pdf
                Reply  Thu May 16 23:31:55 2019, gautam, Update, SUS, ETMY suspension characterization 6x
                   Reply  Fri May 17 17:01:08 2019, gautam, Update, SUS, ETMY suspension characterization etmy_oplevs.pdf
          Reply  Mon May 20 11:33:46 2019, gautam, Update, SUS, ITMY inspection 
             Reply  Mon May 20 17:12:57 2019, gautam, Update, SUS, ETMY LL adjustment 
                Reply  Mon May 20 22:06:07 2019, gautam, Update, SUS, ITMY also kicked 
                   Reply  Tue May 21 00:15:21 2019, gautam, Update, SUS, Main objectives of vent achieved (?) etmy_oplevs_20190520.pdfETMY_cplxTF.pdfETMY_diagComp.pdfScreen_Shot_2019-05-21_at_12.37.08_AM.png
                      Reply  Tue May 21 21:33:27 2019, gautam, Update, SUS, ETMY HR face cleaned 
                         Reply  Wed May 22 11:53:50 2019, gautam, Update, SUS, ETMY EQ stops backed out ETMY_eigenmodes.pdf
Message ID: 14613     Entry time: Thu May 16 13:07:14 2019     In reply to: 14612     Reply to this: 14615
Author: gautam 
Type: Update 
Category: SUS 
Subject: First contact residue removal 

I  used a pair of tweezers to remove the stray fiber of first contact. As Koji predicted, this was rather dry and so it didn't have the usual elasticity, so while I was able to pull most of it off, there is a small spot remaining on the HR surface of the ETM. We will remove this with a fresh application of a small patch of FC.

I the meantime, I'm curious if this has actually fixed the suspension woes, so yet another round of freeswinging data collection is ongoing. From the first 5 mins, looks positive, I see 4 peaks around 1Hz cool!

The following optics were kicked:
ETMY
Thu May 16 13:06:39 PDT 2019
1242072418

Update 730pm: There are now four well-defined peaks around 1 Hz. Together with the Bounce and Roll modes, that makes six. The peak at 0.92 Hz, which I believe corresponds to the Yaw eigenmode, is significantly lower than the other three. I want to get some info about the input matrix but there was some NDS dropout and large segments of data aren't available using the python nds fetch method, so I am trying again, kicked ETMY at 1828 PDT. It may be that we could benefit from some adjustment of the OSEM positions, the coupling of bounce mode to LL is high. Also the SIDE/POS resonances aren't obviously deconvolved. The stray first contact has to be removed too. But overall I think it was a successful removal, and the suspension characteristics are more in line with what is "expected". 

Attachment 1: etmy_sensors.pdf  13 kB  Uploaded Thu May 16 20:34:00 2019  | Hide | Hide all
etmy_sensors.pdf
Attachment 2: etmy_BRmode.pdf  118 kB  Uploaded Thu May 16 20:39:55 2019  | Hide | Hide all
etmy_BRmode.pdf
ELOG V3.1.3-